OpenStudy (anonymous):

i need help please:

OpenStudy (anonymous):

find x if y=-3 \[\frac{ 2x }{ y }+6=-4\]

OpenStudy (anonymous):

Put the value of y as -3 in the equation.

hero (hero):

\[\frac{2x}{-3} + 6 = -4 \\-\frac{2x}{3} = -4 - 6 \\10 = \frac{2x}{3} \\30 = 2x \\\frac{30}{2} = x \]

OpenStudy (hba):

@Hero Forgot the rules ?

OpenStudy (anonymous):

\[\frac{2x}{-3} + 6 = - 4\]

hero (hero):

What are you talking about @hba

OpenStudy (hba):

@Hero You should let the asker participate in the solution rather than giving direct answer or providing the whole solution.

hero (hero):

I left it up to her to divide the last part. 30/2

OpenStudy (hba):

@Hero Well,you should let her do more of that lol. Anyone can divide 30/2 .

hero (hero):

I was showing the process of how to get from the problem to the solution. The student needs to see that. I don't mind showing a student how to do something first time around.

hero (hero):

If she posts a similar problem, then I will challenge her to do it on her own.

hero (hero):

That's the way I do it bro.

hero (hero):

Nothing wrong with it.

OpenStudy (hba):

@Hero Well you could have shown an example and let her do this on her own.

hero (hero):

That is as good as giving an example. Nothing wrong with showing the steps to solve the first problem. Do not challenge me bro. I've been here WAY LONGER THAN YOU.

OpenStudy (anonymous):

im only reviewing my homework.

hero (hero):

@hba, since you have blocked me, I don't want you communicating with me. Your presence insults my intelligence.

OpenStudy (hba):

Because OS is not an answering machine,we do not want them to cheat on their tests and do their homework for them. @Hero Being for a long time does not prove that you will not follow the rules.

OpenStudy (hba):

@Hero Blocked you ?

hero (hero):

Unblock me, then we can discuss this. It is rude to have conversations like this out in the open. Yes, I cannot send you a message.

OpenStudy (hba):

Well let me fan you.

OpenStudy (anonymous):

he never gave me the answer i solved the last part on my own

OpenStudy (aravindg):

exactly @Hero is one of the most experienced and matured users here :)

OpenStudy (hba):

@AravindG Agreed.